OpenStudy (anonymous):

Roland practiced the saxophone for 46 minutes on Monday, 53 minutes on Tuesday, 78 minutes on Wednesday and 60 minutes on Thursday. How many hours and leftover minutes did he practice during these 4 days?

OpenStudy (anonymous):

you add up all the total amounts and then you see how much there is find some hours then get the left over minutes=anwser

OpenStudy (anonymous):

i got 237

OpenStudy (anonymous):

@faithy02

OpenStudy (anonymous):

that 237 is the number of minutes she practiced. if you want to convert that to how many hours, you'll have to divide that by 60

OpenStudy (dumbcow):

60*3 = 180 60*4 = 240 (too much) 237 - 180 = 57 --> 3 hours 57 min

OpenStudy (anonymous):

Thx!!!!!! soooo much

OpenStudy (dumbcow):

your welcome